If 4.3g of mercury(ii) oxide is thermally decomposed to mercury and oxygen , what maw of mercury will be produced. How many molecules will be produced?


(Hg=208,O=16,1 mole=6.02 x 10^23)


1
Expert's answer
2021-09-23T02:34:02-0400

Solution:

(Hg = 200, O=16, 1 mole = 6.02×1023)


Balanced chemical equation:

2HgO(s) ⟶ 2Hg(l) + O2(g)


Calculate the moles of HgO:

The molar mass of HgO is (200 + 16) = 216 g/mol

Hence,

(4.3 g HgO) × (1 mol HgO / 216 g HgO) = 0.0199 mol HgO = 0.02 mol HgO


According to stoichiometry:

2 mol of HgO produces 2 mol of Hg

Thus, 0.02 mol of HgO produces:

(0.02 mol HgO) × (2 mol Hg / 2 mol HgO) = 0.02 mol Hg


Calculate the mass og Hg:

The molar mass of Hg is 200 g/mol

Hence,

(0.02 mol Hg) × (200 g Hg / 1 mol Hg) = 4.0 g Hg


Calculate the number of molecules:

(0.02 mol Hg) × (6.02×1023 molecules / 1 mol) = 1.2×1022 Hg molecules


Answers:

4.0 grams of mercury (Hg) will be produced.

1.2×1022 Hg molecules will be produced.